# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v2

@@ 4.5 LOLA Variants and Extensions
Alternative rewards: In the Upworthy setting, the measure of reward is clicks. However, click-based metrics have raised concerns about promoting clickbait headlines, which may reduce user engagement or spread negativity. As a result, the news and media industry is now evolving to focus more on content quality and humanity, although clicks still remain an important measure; see (Upworthy2023) for a more detailed discussion. The LOLA framework can be easily adapted to alternative reward measures that align with new business goals, e.g., time spent on the platform during a session, or time spent on the article (and not just clicks on the headline).
@@ Appendix A User Survey
Detailed information on the surveys and the results from both surveys are shown in Table A1. Note that the total responses received is less than the number of respondents ×$\times 20$ because some students did not finish the survey or skipped some comparisons. For example, for the first survey, the total number of responses 3,809 is less than $202\times 20=4,040$. To calculate the accuracy metric, we only focus on the significant pairs, and in the first survey, we discard the uncertain answer “Both headlines seem equally good” for a relatively fair comparison between the two surveys. “Accuracy” measure here is consistent with the LLM tests. We treat the headline with a higher CTR as the correct answer and calculate the proportion of correct answers as accuracy. By doing so, one should expect 50% accuracy by random guess since
Detailed information and survey results are documented in Table A1. Some students did not complete the survey or skipped some comparisons, resulting in 3,809 responses out of a potential 4,040 ($202\times 20$). To calculate the accuracy metric, we focused solely on the significant pairs, and in the first survey, we excluded the uncertain answer “Both headlines seem equally good” to facilitate a fair comparison between the two surveys. Accuracy here aligns with the LLM tests. We regarded the headline with the higher CTR as the correct answer and calculated the proportion of correct responses as accuracy. As such, we should expect a 50% accuracy rate from random guessing.
@@ B.1 Pairwise Comparison Between GPT Prompts
To determine whether one prompt is significantly better than the other, we record a binary indicator of the correct/incorrect answer for each prompt on each sample and then conduct a paired t-test on 1,000 correct/incorrect answer indicators to see whether the answers from one prompt are significantly more accurate than the other. The mean differences and p values are documented in Figure A1 and Figure A2. The random guess model simply predicts the first headline wins with a half probability. The nomenclature for GPT prompts in these two figures follows a specific format. The first component is the name of the GPT model, the second component indicates the number of demonstrations, and the third component specifies whether the outputs in the demonstrations are flipped. A value of 0 indicates the correct label, while a value of 1 indicates the wrong label. For example, GPT-4-0-0 means zero-shot prompting using GPT-4, and GPT-3.5-5-1 means in-context learning with 5 demonstrations and wrong labels using GPT-3.5.
@@
Appendix C Appendix to Embedding-based Classification Approaches
C.1 Implementation Details of MLP Classification Models
@@
Hyperparameters of the network structure. We employ a fully connected neural network with ReLU activation functions for the hidden layers and a sigmoid activation function for the output layer. This configuration is widely used in practice. To determine the number of layers and the hidden nodes of each layer, we choose the number of layers within the range of $\{1,2,3,4,5\}$, and set the number of nodes in the first fully connected layer from the set $\{128,256,512,1024,2048\}$. Besides, we adjust the previous-post ratio, which is the proportion of nodes in each layer relative to those in the subsequent layer. This ratio here is in $\{2,4,8\}$. Note that the previous-post ratio becomes irrelevant to the results when there’s only one hidden layer. This is because the final output layer has only one dimension for ranking the most attractive headline. We incorporate the dropout after each hidden layer to mitigate the over-fitting. Dropout layers randomly deactivate neurons based on a specified dropout rate, while scaling up the remaining neurons to maintain the same mean of the layer. In this work, we set the dropout rate in $\{0.3,0.5\}$. All hyperparameters are automatically selected through hyperparameter optimization, which is introduced below.
@@
Hyperparameters of the training procedure. The MLP is implemented in PyTorch, with data split into 70% for training, 10% for validation, and 20% for testing. To prevent the network from memorizing the order of samples, we shuffle the training set at the start of each epoch. Determining the appropriate batch size is crucial, as large batch sizes may lead to convergence to local optima, while small batch sizes may hinder convergence. We balance accuracy and training speed by setting the batch size to 10 and the learning rate to $0.01$. Additionally, to mitigate over-fitting, we apply $L^{2}$ regularization, with the regularization weight chosen from $\{1e-6,1e-4,1e-2\}$. We use binary cross-entropy as the loss function and employ stochastic gradient descent for training. We implement early stopping by monitoring the validation set loss, halting the training process if the reduction in loss over 60 epochs falls below a tolerance threshold of $10^{-4}$.
Hyperparameter optimization. We optimize the parameters mentioned above through Bayesian Optimization (snoek2012practical). This Bayesian technique uses Bayesian inference to construct a probabilistic model of the objective function, where the parameters are updated iteratively with new observations. This iterative process balances exploration of the parameter space to discover promising regions and exploitation of known high-performing areas. In this work, we consider the accuracy of the validation set as the objective and optimize hyperparameters selected from the specified parameter space, as delineated by the ranges mentioned previously. We choose the Gaussian process as the estimator with the automatic optimizer, which is commonly adopted by the community. To ensure a judicious trade-off between efficacy and efficiency, we repeat the optimization process 50 times for input data of 256 dimensions and 30 times for data of 3,072 dimensions.
C.2 Effect of Headline Similarity on Performance
We now examine the conditions under which these approaches fail. , Specifically, we investigate whether headline similarity affects the performance of embedding-based classification models. We first calculate the cosine similarity based on the text embeddings for each pair of headlines. This metric captures the similarity between two vectors in inner product space and is defined as the dot product of the vectors divided by the product of their lengths, i.e., ${e_{1}\cdot e_{2}}/(\|e_{1}\|\|e_{2}\|)$, where $e_{1}\cdot e_{2}$ is the dot product of the embedding vectors, and $\|e_{1}\|$ and $\|e_{2}\|$ are the their Euclidean norms. Figure 3(a) shows the distribution of cosine similarity across all headline pairs, whereas Figure 3(b) shows the distributions of cosine similarity separately for two cases: (1) cases where our embedding-based classification correctly predicts which headline is catchier, and (2) cases where our embedding-based classification incorrectly predicts which headline is catchier. We observe a rightward shift in the distribution of incorrect predictions (compared to the distribution of correct predictions). This suggests that incorrect predictions are more common when the headlines are very similar to each other (i.e., higher cosine similarity). There are two possible reasons for this. First, in cases where the headline pairs are very similar, there is a higher chance of Type 1 error (in the t-tests based on our A/B test data), i.e., the headlines may not be significantly different in fact. Second, text embeddings may not be able to capture the differences between two headlines when they are similar, i.e., the embeddings may also be very similar. As a result, it is hard for the classification model to differentiate between them. [^19] Note that it is unlikely that this phenomenon is due to problems/issues with the classification model because both Logit and MLP achieve nearly identical performance.
(a) Cosine similarity distribution over all pairs in training and test sets
(b) Cosine similarity over incorrect/correct predicted test pairs
Figure A3: The left panel shows the distribution of cosine similarity for all pairs. The right panel shows the distributions of cosine similarity for both incorrect and correct predictions using the logit classification model with OpenAI-3072E embeddings.
@@ C.3 Effect of Training and Test Data Size on Performance
(a) Performance of OpenAI-256E
(b) Performance of OpenAI-3072E
@@
Figure A4: Performance of Logit classification models with OpenAI embeddings under varying training dataset sizes and different embedding sizes.
We now conduct experiments varying the proportion of data used for training vs. test. We use the Logit model for these experiments since it offers the best performance and is easy to train. The results from this exercise are shown in Figure A4. As we can see, when we use the 256-dimensional embeddings, approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high and stable performance. However, when we use the 3072E-dimensional embedding, there is a clear over-fitting [^20] Over-fitting refers to the phenomenon that the accuracy on the training set is consistently higher than the accuracy on the test set. It occurs when a machine learning model learns the details and noise in the training data to the extent that it negatively impacts the model’s performance on new, unseen data. This often happens because the model becomes too complex or because the training set is too small, causing the model to memorize specific examples and noise rather than general patterns. issue even if we use the simplest Logit model and 95% of the data for training. This suggests that for larger embedding vectors, we need much larger training datasets. [^21] In principle, overfitting can also be due to model complexity. However, this is unlikely to be the case in our setting since we are using a simple Logit model with linear features.
C.4 Dimensionality Reduction and Regularization
In this section, we examine whether employing dimensionality reduction techniques and/or regularization can improve performance and reduce overfitting.
C.4.1 PCA for Embeddings
One can observe that OpenAI-256E in Figure A4(a) does not suffer from the over-fitting in contrast to OpenAI-3072E as shown in Figure A4(b). Thus, one straightforward method one can try to attack the over-fitting issue is reducing the dimension of embeddings using PCA or directly using lower-dimensional OpenAI embeddings. However, as we have discussed that OpenAI-256E has a worse performance than OpenAI-3072E regarding the accuracy in the testing dataset, we turn to PCA to see if PCA-reduced embeddings can improve the performance by solving the over-fitting issue.
Specifically, we first use PCA to reduce 3072-dimensional OpenAI embeddings. Then, we concatenate two reduced embeddings from healing pairs and use Logit for the downstream classification task. The result is displayed in Figure A5, which shows that as the dimension is reduced, although there is a smaller accuracy gap between the training and test set, the accuracy on both datasets decreases, indicating that PCA does not work for our task, which prefers higher testing accuracy.
Figure A5: Accuracy of the Logit classification model with varying PCA-reduced OpenAI embeddings as inputs.
C.4.2 $L^{1}$, $L^{2}$ Regularization
[…]
Figure A6: Accuracy of the Logit classification model with varying $L^{1}$ (left) and $L^{2}$ (right) regularization weights. X-axis is the regularization weight defined as $1/\texttt{C}$.
In sum, we find that neither PCA nor regularization improves the performance accuracy on the test data, which suggests that over-fitting is unlikely to be due to learning noise in the training data. Our analysis suggests that we need more A/B tests to learn the relationship between a headline’s text embeddings and its attractiveness effectively. In summary, these comparisons suggest that data samples from Upworthy are insufficient to unlock the potential of LLMs fully. Nevertheless, diverting more and more traffic to A/B tests is costly, and this observation is one of the motivations for our approach later in §$\S$4.
C.5 Potential Threats and Robustness Checks
@@
Next, we discuss two potential threats to the inference from our analysis here and our robustness checks to alleviate concerns around these threats. First, one may be concerned as to whether Upworthy’s data has been used as part of OpenAI’s training corpus – which could potentially invalidate our analysis. However, this is unlikely for several reasons. First, it is improbable that OpenAI would have specifically downloaded these Upworthy CSV files (Comma-Separated Values files) dataset from a subfolder in Upworthy Research Archive. Second, the original data only contains impressions and clicks for each headline, as shown in Table 1, rather than paired headlines with winner information. Therefore, during the training of OpenAI models, it is unlikely they were exposed to the winner information since the training objective is to predict the next word. Third, we conducted an experiment to determine if this dataset was indeed a part of OpenAI’s corpus. We used embeddings from two headlines with significantly different impressions (possibly from different tests) for the downstream task of predicting which headline received higher impressions. If the embedding model was trained on this corpus, its embeddings should reflect the number of impressions and perform well on this task. However, the observed accuracy was approximately 50%, indicating that the embeddings do not contain such information, confirming our assumption. Nevertheless, if a firm is concerned that its data is a part of public repositories or is concerned about privacy issues around submitting headlines to OpenAI to obtain embeddings, then instead of using embedding-based approaches, it can rely on fine-tuning based approaches; see §$\S$3.3 for details.
@@
Second, one may be concerned that readers’ preferences changed over time and that the current sample splitting method fails to take into account when the headline was created, making our classification task invalid. To address this concern, we conducted a robustness check by splitting the dataset based on the creation time of the headlines (e.g., the initial 80% of the data, in chronological order, as the training set, and the subsequent 20% as the test set). We found its accuracy was 81.50% (close to the best embedding accuracy of 82.50%), indicating that time shift preference is not a concern in our data.
@@ D.1 Transformer and Llama Architecture
It is worth noting that LLMs do not necessarily operate in a conversational manner, nor do they always use language as their output. As we will demonstrate in §$\S$3.3.2, we employ the Llama model as a classifier, where the output is the distribution of predicted classes. In applications like conversational AI, exemplified by GPT, the contextualized embeddings from the last token of the last layer of the Transformer are fed into a linear layer followed by a softmax function, the configuration of which transforms these embeddings into a probability distribution over the vocabulary, enabling the model to predict the next token in the sequence.
D.2 Additional LoRA Models
D.2.1 Incorporating Insignificant Pairs for LoRA Fine-tuning
Similar to the embedding-based logit model, we have experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay in an attempt to mitigate over-fitting. However, as we observed, these modifications generally reduce performance on the training set without yielding significant improvements on the test set.
In the following, we test whether incorporating more data, including insignificant pairs, can help improve the performance. Notice that by incorporating more significant pairs, we indeed observed improvement in test accuracy (see Figure A4), the observation of which holds for both embedding-based models and LoRA fine-tuned models.
Here, we only present the results from fine-tuning Llama-3-8b using different sizes of training data, and the result from the embedding-based model remains the same. Recall that in §$\S$3.3.3, we use only significant pairs (with p-values smaller than or equal to 0.05, which includes 39,158 pairs) and split them into training and test sets. We continue to use the same significant test set for performance evaluation. However, for the training set, we obtain the data in the following manner: first, we remove the fixed test set from all pairs, leaving a total of 134,941 pairs. Then, given a ratio (e.g., 60%), we select these remaining pairs’ first 60% (ranked by ascending p-value) as the training set, see Figure A8 for a diagram of our dataset construction.
Figure A8: Dataset Construction by Incorporating Both Significant and Insignificant Pairs in Llama-3-8b LoRA Fine-tuning
We set the ratio to 20%, 40%, 60%, 80%, and 100%. For each ratio and size of the training set, we finetune the Llama-3-8b model and evaluate the performance on the fixed test set. Except for the training set, all other training configurations are the same as in §$\S$3.3.2. The performance is reported in Figure A9. The accuracy of the training set decreases as the ratio gets larger since we are incorporating more insignificant pairs. The accuracy on the test set maintains around 83%, which indicates that incorporating more training data does not necessarily lead to better performance, at least for the current dataset. Notice that the decreasing trend of accuracy on the training set in Figure A4 is less steep because the pairs in that training set are all significant. However, the main goal here is to confirm that more insignificant pairs cannot improve the performance for either LoRA or the embedding method.
Figure A9: Performance of fine-tuning Llama-3-8b under varying training dataset sizes.
D.2.2 DistilBERT LoRA Fine-Tuning
We now discuss One difference between DistilBERT and Llama-3 is that DistilBERT is a bidirectional Transformer model, which means that its attention mechanism allows the model to learn the context from both directions, while Llama-3 is a unidirectional model. An illustration of the structure of the DistilBERT model combined with a classification head is shown in Figure A10. By default, DistilBERT can handle two headlines together as input. The format of the input is: [CLS] Headline 1 [EOS] Headline 2 [EOS]. The [CLS] token is a special token used to get the output of the model for classification, and the output of the last Transformer layer corresponding to this token is used as the input to the classification head. The [EOS] token is used to separate the two sentences.

### v3

@@ 4.5 LOLA Variants and Extensions
Alternative rewards: In the Upworthy setting, the measure of reward is clicks. However, click-based metrics have raised concerns about promoting clickbait headlines, which may reduce user engagement or spread negativity. As a result, the news and media industry is now evolving to focus more on content quality and longer-term user engagement and retention metrics, although clicks still remain an important measure; see (Upworthy2023) for a more detailed discussion. The LOLA framework can be easily adapted to alternative reward measures that align with new business goals, e.g., time spent on the platform during a session, or time spent on the article (and not just clicks on the headline).
@@ Appendix A User Survey
Detailed information and survey results are documented in Table A1. Some students did not complete the survey or skipped some comparisons, resulting in 3,809 responses out of a potential 4,040 ($202\times 20$). To calculate the accuracy metric, we focused solely on the significant pairs, and in the first survey, we excluded the uncertain answer “Both headlines seem equally good” to facilitate a fair comparison between the two surveys. Accuracy here aligns with the LLM tests. We regarded the headline with the higher CTR as the correct answer and calculated the proportion of correct responses as accuracy. As such, we should expect a 50% accuracy rate from random guessing in this paired comparison.
@@ B.1 Pairwise Comparison Between GPT Prompts
To determine whether one prompt is significantly better than the other, we record a binary indicator of the correct/incorrect answer for each prompt on each sample and then conduct a paired t-test on correct/incorrect answer indicators to see whether the answers from one prompt are significantly more accurate than the other. The mean differences and p values are documented in Figure A1 and Figure A2. The nomenclature for GPT prompts in these two figures follows a specific format. The first component is the name of the GPT model, the second component indicates the number of demonstrations, and the third component specifies whether the outputs in the demonstrations are correct or not. A value of 0 indicates the correct label, while a value of 1 indicates the wrong label. For example, GPT-4-0-0 means zero-shot prompting using GPT-4, and GPT-3.5-5-1 means in-context learning with 5 demonstrations and wrong labels using GPT-3.5.
@@
Appendix C Appendix to Embedding-based Approaches
C.1 Implementation Details
We start by introducing three embedding models used:
  • $\bullet$
    OpenAI embeddings: At the time of this study, the latest and best embedding model available from OpenAI was text-embedding-3-large. These embeddings are obtained using the OpenAI embedding API in Python. This embedding model generates embeddings with up to 3,072 dimensions and supports a maximum token limit of 8,191, with its knowledge base extending up to September 2021. A notable feature of this model is its flexibility – it allows users to tailor the embedding dimension to their specific needs. [^16] Other models capable of producing text embeddings, like BERT (devlin2018bert) and Sentence-BERT (reimers2019sentence), have fixed embedding vector dimensions that remain constant once the model is trained. These dimensions cannot be changed or resized for different applications. However, OpenAI’s embedding model uses a technique that, while producing outputs of fixed length, encodes information at different granularities (kusupati2022matryoshka). This allows a single embedding to adapt to the computational constraints of downstream tasks. More specifically, downstream tasks can use only the first several dimensions of the original vector, and the information retained decreases smoothly as a shorter part of the vector is used. In our experiments, we consider both 3,072- and 256-dimensional embeddings to examine whether the size of the embedding vector size affects the performance of the downstream CTR prediction models. [^17] A higher dimension does not necessarily mean better performance. While a higher dimension can retain more information for each sample, it also means that the samples are spread more sparsely in a higher-dimensional space given a limited number of training samples. This sparsity can increase the possibility of over-fitting.
  • $\bullet$
    Llama-3-8b: Meta’s Llama-3 (metallama3) is regarded as one of the most advanced open-source LLMs to date (llama3modelcard). We obtain embeddings from Llama-3 with 8 billion parameters (Llama-3-8b). We fine-tune this exact model using LoRA in §$\S$3.3; therefore, using the embeddings from this model allows for a direct comparison of fine-tuning vs. embedding-based methods (see §$\S$3.4 for comparisons). Llama-3-8b does not produce embeddings explicitly designed to represent the semantic meaning of entire input sequences. To address this issue, we apply the technique described in (behnamghader2024llm2vec), which transforms the Llama-3-8b model into a more suitable text embedding model, outputting embeddings with a dimension of 4,096.
  • $\bullet$
  • $\bullet$
    BERT embeddings: BERT (Bidirectional Encoder Representations from Transformers), introduced by devlin2018bert, is a seminal model in natural language processing that generates context-sensitive embeddings for text. Unlike earlier models that produce static word embeddings, BERT captures the nuances of word meanings based on their context within a sentence, resulting in more accurate representations. The standard BERT model produces embeddings with 768 dimensions, which are fixed once the model is trained. This fixed dimensionality can be a limitation when adapting to tasks with varying computational constraints. Despite this, BERT’s ability to understand context has led to its widespread adoption in various NLP applications, including sentiment analysis, question answering, and named entity recognition.
To assess the impact of model complexity on performance, we consider using a simple linear model and a multilayer perceptron (MLP) model. MLP is a powerful neural network architecture consisting of multiple linear and non-linear hidden layers that allow the model to learn complex patterns in the data. We use a fully connected neural network with ReLU activation for hidden layers, with dropout layers added to prevent over-fitting. We refer interested readers to Chapter 5 in (zhang2021dive) for more details. We train the MLP using the stochastic gradient descent algorithm to minimize binary cross-entropy and perform hyper-parameter tuning with Bayesian Optimization to determine the optimal configuration, with the training process and the architecture details below.
@@ C.1 Implementation Details
Hyperparameters of the network structure. We employ a fully connected neural network with ReLU activation functions for the hidden layers, a configuration widely used in practice. The network consists of two hidden layers, each with 128 nodes. To mitigate overfitting, we incorporate dropout after each hidden layer. Dropout layers randomly deactivate neurons based on a specified dropout rate while scaling up the remaining neurons to maintain the same mean of the layer. We set the dropout rate to 0.5. We did not further test other hyperparameter settings, as this configuration already delivers sufficiently good performance. However, more careful hyperparameter optimization could potentially improve results.
@@
Training Procedure Hyperparameters. The MLP is implemented in PyTorch, with the data split into 70% for training, 10% for validation, and 20% for testing, as introduced in §$\S$3 . To prevent the network from memorizing the order of samples, we shuffle the training set at the start of each epoch. Determining the appropriate batch size is crucial, as large batch sizes can lead to convergence to local optima, while small batch sizes may slow down convergence. We balance accuracy and training speed by setting the batch size to 10 and the learning rate to $0.01$. The loss is calculated as the mean squared error between the real CTR and the predicted CTR, and we use stochastic gradient descent for training.
C.2 Performance Comparison of Different Embeddings
Embedding Model CTR Prediction Accuracy on the test set
OpenAI-256E Linear 46.28%
MLP 44.38%
OpenAI-3072E Linear 43.06%
MLP 45.17%
Llama-3-8b-4096E Linear 42.78%
MLP 36.03%
Word2Vec-256E Linear 41.07%
MLP 36.96%
Word2Vec-3072E Linear 40.27%
MLP 37.27%
BERT Linear 42.32%
MLP 35.89%
Table A2: The performance of different embedding and CTR prediction models on test data. All models use the same training data, and accuracy is evaluated on the test data. Note that the Linear model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents. Therefore, we rerun MLP training three times with different random seeds and report the average accuracy from three reruns.
The comparison of different embeddings and CTR prediction models are listed in Table A2. We observe that OpenAI embeddings outperform Llama-3-8b, Word2Vec, and BERT-based models. This is because OpenAI’s LLMs are state-of-the-art models that naturally perform better than the other embeddings. Additionally, the performance of embeddings from the open-source Llama-3-8b falls between that of the proprietary OpenAI model and Word2Vec/BERT, which does not use the transformer architecture.
@@ C.3 Effect of Training and Test Data Size on Performance
In this analysis, we focus on OpenAI-256E (as it provides the best performance) and use linear regression as our prediction model, since the MLP does not significantly outperform the linear model on OpenAI embeddings and is easier to train.
@@
Figure A3: Performance of CTR prediction model with OpenAI-256E under varying training dataset sizes
To assess the effect of training sample size on accuracy, we vary the training data size and measure performance. Specifically, we set a training size ratio, randomly select a subset of articles based on this ratio from the full training set, and use this subset to train the model. As shown in Figure A3, increasing the amount of training data leads to a gradual improvement in test accuracy. For example, increasing the training size from 20% of the total dataset (about 30% of the training set) to 70% of the total dataset (100% of the training set) raises accuracy from 44% to 46% on the test data. This observation also indicates that, achieving a significant performance boost in LLMs would require a much larger dataset than is currently available.
C.4 Potential Threats and Robustness Checks
Note that in the following two analyses of threats, we still focus on OpenAI-256E with linear regression as our prediction model as it provides the best performance.
@@ C.4 Potential Threats and Robustness Checks
Next, we discuss two potential threats to the inference from our analysis here and our robustness checks to alleviate concerns around these threats. First, one may be concerned about whether Upworthy’s data has been used as part of OpenAI’s training corpus, which could potentially invalidate our analysis. However, this is unlikely for several reasons. First, it is improbable that OpenAI would have specifically downloaded these Upworthy CSV files (Comma-Separated Values files) dataset from a subfolder in Upworthy Research Archive. Second, the original data only contains impressions and clicks for each headline, as shown in Table 1, rather than CTR information. Therefore, during the training of OpenAI models, it is unlikely they were exposed to the CTR information since the training objective is to predict the next word. Third, we further conducted an experiment as the robustness check to determine if this dataset was indeed a part of OpenAI’s corpus. We used embeddings from two headlines with significantly different impressions (possibly from different tests) for the downstream task of predicting which headline received higher impressions. If the embedding model was trained on this corpus, its embeddings should reflect the number of impressions and perform well on this task. However, the observed accuracy was approximately 50%, indicating that the embeddings do not even contain such impression information, confirming our assumption.
@@
Second, one may be concerned that readers’ preferences changed over time and that the current sample splitting method fails to take into account when the headlines were created, potentially making our CTR prediction model invalid if users’ preferences shift. To address this concern, we conducted a robustness check by splitting the dataset based on the creation time of the headlines. Specifically, we used the initial 70% of the data, in chronological order, as the training set, and the latest 20% as the test set (the remaining 10% was reserved for algorithm hyperparameter fine-tuning). We found that the accuracy was 46.25% on the test data, which is close to the embedding performance on the randomly split training and test dataset, with an accuracy of 46.28%, as given in Table 5. This indicates that time-based preference shifts are not a concern in our data.
@@ D.1 Transformer and Llama Architecture
It is worth noting that LLMs do not necessarily operate in a conversational manner, nor do they always use language as their output. As we will demonstrate in the following, we employ the Llama model as a CTR prediction model. In applications like conversational AI, exemplified by GPT, the contextualized embeddings from the last token of the last layer of the Transformer are fed into a linear layer followed by a softmax function, the configuration of which transforms these embeddings into a probability distribution over the vocabulary, enabling the model to predict the next token in the sequence.
D.2 LoRA Fine-Tuning Technique
We illustrate LoRA in Figure A5. We use $W$ to represent any pre-trained weight matrix, which can be either $W_{Q}$, $W_{V}$, or $W_{K}$. Specifically, $W_{Q}$ is the weight matrix for the query vectors, $W_{K}$ is for the key vectors, and $W_{V}$ is for the value vectors in the attention mechanism of the Transformer architecture. These matrices store most of the information and knowledge learned by LLM, and are crucial for calculating the attention scores and subsequently determining the relevance of different tokens in the input sequence. For more technical background, we refer to Web Appendix §$\S$§D.1. We use $d\times d_{W}$ to represent the dimension of the weight matrix and $r$ to represent the rank of the low-rank matrices. We denote the modified weight matrices after fine-tuning as $W^{\prime}$, which can be further expressed as:
Figure A5: Illustration of LoRA Fine-Tuning.
D.3 Implementation Details
We use exactly the same training and test sets as in §$\S$3.2, but the inputs here are the original headlines with the text information. We add a regression head to the original Llama-3-8b base model, which includes 32 Transformer layers, as depicted in Figure A6. Llama-3-8b employs a causal attention mask, where the attention is unidirectional—each token can only attend to preceding tokens in the sequence. Thus, the last token in the sequence, being informed by all prior tokens, is used for predicting the CTR. This embedding vector of the last token (which is 4096-dimensional), output by the last Transformer layer, is processed through a linear layer producing a one-dimensional scalar that represents the predicted CTR.